4. Vanguard Office Supplies is a nationwide retail chain that offers office supplies and office furniture. Company management has decided that, from both a competitive and a cost-cutting standpoint, Vanguard should offer its own private-label brands for products like student notebooks, fillers, ledgers and journals, bond and linen paper, and other products. To accomplish this objective, Vanguard is considering the purchase of Omega Paper, a manufacturer of paper products and notebooks. A five-year income forecast for Omega is given, along with other pertinent information. Vanguard plans to keep Omega's debt-equity ratio at its current level [Use spreadsheet Valuation Workbook provided]

Five-Year Forecast ($ MM)

Year 1

Year 2

Year 3

Year 4

Year 5

Revenue

957.9

972.2

986.8

1,001.6

1,016.6

Cost of goods sold

862.1

875.0

888.1

901.4

915.0

SG&A

67.0

68.1

69.1

70.1

71.2

EBIT

28.7

29.2

29.6

30.0

30.5

Interest expense

12.5

13.4

13.8

14.1

14.4

EBT

16.2

15.7

15.8

16.0

16.1

Taxes

5.5

5.3

5.4

5.4

5.5

Net Income

10.7

10.4

10.5

10.5

10.6

Other projections:

Change in net working capital

(5.7)

(5.8)

(5.9)

(6.0)

(6.1)

Capital expenditures

28.7

29.2

80.3

62.2

30.5

Depreciation

52.7

53.5

54.3

55.1

55.9

Tax rate                                                                    34.0%

Book value of debt ($MM)

192.3

Book value of equity ($MM)

364.9

Share price ($)

10.5

Beta

0.76

Shares outstanding (MM)

51.5

Interest rate on company debt

6.5%

Current bond yields on similarly rated companies

6.1%

10-year T-bond rate

5.0%

Inflation

2.5%

Market risk premium

6.0%

Calculate the equity value in total and per share of Omega using the discounted cash flow method (show all your calculations for credit).
Answer #1

SOLUTION:-

Particulars Year-1 2 3 4 5
EBIT 28.70 29.20 29.60 30 30.50
Less:Tax @ 34% 9.80 9.90 10.10 10.20 10.40
NOPAT 18.90 19.20 19.50 19.80 20.10
Add:-Deprecaition 52.70 53.50 54.30 55.10 55.90
Less:Capital Exp 28.70 29.20 80.30 62.20 30.50
Add:- Negatvive Change in net W.Capital 5.7 5.8 5.9 6 6.1
FCFF 48.60 49.30 -0.6 18.70 51.60

..

After tax cost of debt = 6.1*0.66 = 4.03%

Cost of Equity = 5 + 6*0.76 = 9.56%

..

Wacc = 7.65%

..

Value of Firm = $811.33

VALUE OF EQUITY = Value of firm - value of debt

= 811.33 - 192.30

= $619.03 m

..

Per share value = $12.02